# What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality

### v2

@@ Abstract
Mixed Reality (MR) enables rich, embodied collaboration, yet it is uncertain if sensor and system-logged behavioral signals capture how users experience that collaboration. This disconnect stems from a fundamental gap: behavioral signals are observable and continuous, while collaboration is interpreted subjectively, and shaped by internal states like presence, cognitive availability, and social awareness. Our core insight is that sensor signals serve as observable manifestations of subjective experiences in MR collaboration, and they can be captured through sensor data such as shared gaze, speech, spatial movement, and other system-logged performance metrics. We propose the Sensor-to-Subjective (S2S) Mapping Framework, a conceptual model that links observable interaction patterns to users’ subjective perceptions of collaboration and internal cognitive states through sensor-based indicators and task performance metrics. To validate this model, we conducted a study with 48 participants across 12 MR groups engaged in a collaborative image-sorting task. Our findings show a correlation between sensed behavior and perceived collaboration, particularly through shared attention and proximity.
@@
Modern MR systems generate extensive behavioral data through system logs and embedded sensors, capturing where users look [47], how they move [34], and when they speak [52], yet these signals often provide only limited insights into users’ internal states [73]. Prior work shows that speech, gaze, and turn-taking are tightly connected to the internal perception of collaboration [11]. However, two core limitations remain: (1) it is unclear whether sensor data meaningfully reflects how participants perceive collaboration, or if key experiential dimensions go undetected. (2) Existing studies demonstrate specific bi-modal correlation mappings such as gaze to presence [4], and physiological synchrony to co-presence [9], but stop short of offering a holistic objective-to-subjective mapping of collaboration.
We address these gaps through a systematic, multimodal Sensor-to-Subjective (S2S) mapping framework, applied in a multi-perspective study of collaborative MR involving 48 participants across 12 teams. First, we analyze sensor-derived indicators of group interaction, including gaze-based shared attention, conversation dynamics through audio, and spatial proximity, which can be captured through embedded sensing in MR headsets [73, 70]. Second, we examine task performance using system logs, measuring timing, interaction patterns, and decision changes throughout the collaboration. Finally, we assess subjective experience through post-exposure questionnaires that capture participants’ perceptions of group dynamics as well as their individual sense of presence and cognitive effort.
Together, these perspectives allow us to ask: Do observable group interaction patterns, as captured by MR headsets, reflect how participants experience collaboration? We formalize this inquiry through three research questions:
@@
Do objective sensor-based indicators of group behavior align with participants’ subjective perceptions of collaboration in MR?
@@
What insights can we derive about group behavior in MR from objective sensing and individual subjective experiences?
@@ 1.1 Group Behavior in Immersive Environments
Understanding collaborative group behavior has evolved from traditional settings to MR environments, where users interact in novel ways across co-located, remote, and hybrid configurations. Group behavior, all forms of interaction and activity within a group [22], and collaboration, defined as the intentional, coordinated effort among members to work together toward shared goals [22], takes distinctive forms in MR contexts due to unique factors including free movement [44], embodied gesture [27], co-presence [55], and shared spatial context [43]. Research has examined MR collaboration through interface design [7], task coordination [71], communication patterns [3], and trust development [6], yet it generally relies on external observations or outcomes rather than internal experiences of the participants. This creates a critical knowledge gap on how participants’ internal experiences align with the observable behaviors logged by MR systems.
@@ 1.2 Sensor-Based Indicators of Group Behavior
Sensor-based methods across ubiquitous and immersive computing characterize group behavior through multiple data streams (gaze, voice, motion, location) without relying on external observation or self-reporting. Remote collaboration systems, such as TeamSense [73] and CoCo [50] monitor nonverbal cues to track group cohesion, while wearable systems assess synchrony and proximity patterns during collaborative tasks [58]. In MR, headset-embedded sensors can track gaze, object manipulation, and gesture alignment, which researchers have shown correspond to collaboration quality [70, 26]. Studies have correlated gaze and gesture sharing with co-presence [4], and linked physiological synchrony and gaze with perceived collaboration quality [9]. Other systems enhance spatial awareness by embedding gaze cues [28], while multimodal integrations of gaze and gesture have been used to improve performance and presence [5]. Additional work shows high-accuracy cognitive load prediction using multimodal sensing [25] and demonstrates that even minimal AR cues can influence team experience and [61]. Conversation-focused research further emphasizes how speech alignment, gesture, gaze, and turn-taking synchronize to create effective collaboration in co-located settings [60, 11]. These findings emphasize that collaboration is fundamentally multimodal, involving a convergence of signals across time.
Despite these insights, existing work remains limited in two ways: (1) isolated pairwise mappings rather than producing a generalizable mapping model and (2) lens-like views often aim to enhance either performance outcomes or structural interaction properties (such as coordination patterns and role distribution) without examining their relationship to participants’ subjective experiences. Prior work has explicitly recognized the need for structured frameworks that integrate multiple sensing modalities and interpret internal collaboration states. Barrett et al. call for multi-method frameworks that integrate sensor modalities to capture internal collaboration states [48], while social signal processing emphasizes the need to fuse multimodal behavioral signals to capture fine-grained group dynamics [61]. Systematic reviews also note the absence of frameworks that link objective data to users’ subjective experiences [17]. Despite these advances, the critical question of whether sensor-based behavioral indicators meaningfully reflect users’ internal perceptions of collaboration remains largely unexplored, particularly in co-located MR tasks using commercial headsets.
@@ 1.3 Subjective Indicators of Group Behavior
Subjective self-report surveys remain central to understanding how people experience collaboration. Standardized tools, such as post-task surveys, presence questionnaires [68, 51], and cognitive load surveys [23], are widely used in the Human-Computer Interaction (HCI) to gauge how users evaluate, interpret, and internalize virtual experiences. Prior work has used these subjective measures to assess collaborative quality in various contexts, including team cohesion in distributed work [73], perceived social presence in remote collaboration [45], and mutual understanding during joint problem-solving [72]. Educational studies use them to evaluate contribution equity and group satisfaction [67], while collaborative design work leverages them to assess idea sharing and creative synergy [41].
Despite rich insights, these subjective measures are rarely systematically linked to sensor-based behavioral data. This disconnect raises key questions: can gaze, speech, and proximity signals in MR environments meaningfully reflect users’ internal sense of collaboration? Do behavioral patterns align with perception, or challenge our assumptions? Answering these questions requires bridging sensor data and subjective experience in collaborative MR.
@@
2 Approach
2.1 Theoretical Foundation
We conceptualize how group behavior can be interpreted by combining interaction traces captured through sensors and system data with subjective user reports.
2.1.1 Subjective Perception of Collaboration
We define collaboration as not just joint actions, but also how individuals interpret, feel, and reflect on their shared experience. In MR, where social cues, task progress, and attention are distributed across a blend of physical and virtual elements, shaped by both individual states and group-level awareness. We treat this as the subjective perception of collaboration and operationalize it through individual-level and group-level dimensions.
Individual-Level Experience. Each user’s interpretation of collaboration is filtered through their sense of presence [54], cognitive availability [56], and personal contribution [14]. These passive background states directly influence how users perceive and make sense of social interaction [45]. At its most foundational, collaboration begins with how individuals experience being part of a group in each moment. In immersive settings, this includes:
Cognitive Load: The mental effort to manage task demands affects users’ ability to perceive coordination signals. A taxed user may miss coordination cues, while a cognitively available one may detect attention shifts or conversational breaks [15]. Thus, we hypothesize that, cognitive load influences not only task performance [30] but also perceived collaboration quality [59].
Contribution Awareness: Users’ sense of how meaningfully others contributed serves as a proxy for perceived participation balance [13]. This acts as a bridge between personal effort and social perception. It links internal judgment with observed peer input and becomes especially salient in unstructured, role-free MR settings.
These individual experiences are not isolated metrics. They are influenced by how users interpret others’ behaviors in real time. They form the perceptual and cognitive foundation on which group-level reflections are built.
Group-Level Reflection. MR system tracks each user individually and users form post-task judgments about shared experience [57], such as whether they attended to the same objects, communicated effectively, and collaborated cohesively. We capture this through four constructs:
@@ 2.1.1 Subjective Perception of Collaboration
Group Collaboration measures users’ overall sense of team effectiveness based on observed behaviors, internal judgments, and implicit comparison to collaborative expectations [19].
These individually collected measures reflect each user’s interpretation of the group, shaped by their state, engagement, and task involvement, forming subjective readings of an inherently social phenomenon.
2.1.2 Objective Perception of Collaboration
While subjective reports capture how collaboration feels, immersive systems log how it unfolds. Although these are individual-level measurements, they are not isolated; they reflect what each user does in relation to others in the group. The same user performing the task alone would not produce the same sensor traces. These sensor-based signals (derived from sensor data and system instrumentation) are shaped by interaction: who a user looks at, how close they are to others, and when they speak or remain silent. Interpreting such distributed traces in group contexts is key to understanding collaboration objectively. We focus on three core indicators supported by prior work [70, 73] and accessible on commodity MR headsets [37, 1, 2]:
Conversation and speech activity have long been used to study group behavior across psychology, anthropology, and HCI [33, 12]. Recent work using ubiquitous sensors underscores the continued importance of conversation [32]. In MR, where interaction blends physical and virtual contexts, speech activity offers valuable insight into collaboration [70]. To this end, logged via headset microphones, we extract speaking time, turn count, and participation variance (non-transcribing content). These metrics, known to reflect dominance and fluency [42], offer insight into how evenly members contribute.
Shared attention, the ability to jointly focus on an object, underpins coordination and social connection [69]. It has been widely studied in collaborative systems [8, 39], often through gaze-based detection[72]. We build on this by using MR headset eye-tracking to detect overlapping gaze on the same virtual object, enabling continuous, time-resolved logging of shared attention [39]. Though not explicitly communicated, such alignment reflects group awareness and may shape perceived collaboration. We aim to assess whether sensed shared attention reflects users’ perception of collaboration.
@@ 2.1.2 Objective Perception of Collaboration
We focus on group-level distributions of proximity and conversation (overall closeness or speech participation) rather than directional relationships (who was closest to whom or who spoke to whom). Our goal is to capture collective engagement, not dominant roles. Prior work shows that proximity itself (regardless of direction) signals social engagement [16], and evenly distributed proximity and turn-taking indicate group cohesion [73]. Moreover, our subjective measures focus on perceived group collaboration, which is more likely influenced by general closeness than specific spatial arrangements.
Beyond sensor signals, we log task-relevant interactions such as how often each user interacts with content, how frequently prior actions are revised, and the number of unique configurations a group explores before reaching a consensus. These metrics reflect participation balance, coordination, and decision negotiation. We also record task completion time as a coarse measure of collaborative success.
2.2 Individual Reflections and Collaboration
@@
A central question in our study is how individual perceptions of collaboration relate to the sensed behavioral traces. Rather than averaging subjective reports into a “single group state”, we examine the distribution of perceptions across the group: Are they aligned or divergent? Do some users feel disengaged or overloaded? Such differences reflect varied interpretations of what collaboration felt like to the user. By linking sensor-based indicators (§2.1.2) with subjective internal perceptions (§2.1.1), we ask: Do shared gaze events map onto users’ awareness of shared attention? Does proximity align with perceived cohesion? Do balanced interactions align with users’ perception that they and their peers contributed meaningfully to the task?
@@
Importantly, the subjective states are not outcomes of the sensor data; they are independent yet complementary accounts of collaboration. Our framework treats them as the interpretive anchor against which we examine the dynamics captured by the system. This allows us to understand what groups did and how collaboration was experienced in the future, yielding insight into how systems might better reflect, support, or adapt to group behavior in MR environments. Together, these indicators form a continuous stream of behavioral data interpreted as isolated features but as observable patterns of interaction that may or may not align with how users interpret collaboration subjectively. Rather than labeling groups as “effective” or “cohesive,” our goal is to assess whether and how the system’s observations reflect how collaboration was actually experienced.
2.3 Sensor-to-Subjective Mapping Framework
We introduce the Sensor-to-Subjective Mapping Framework (S2S) (Figure 1) to structure our investigation. It conceptualizes collaborative group behavior in MR across three perspectives: observable system-level interaction traces (sensor-based indicators), emergent outcomes from interaction (task performance metrics), and user-reported post-task reflections (subjective perceptions). Unlike prior work [4, 9, 48, 60], assuming one view determines another, our framework treats these perspectives as complementary, each capturing different facets of collaboration as described in §1.2.
2.3.1 From Signals to Perception

### v3

# What Sensors See, What People Feel: An Exploratory Study of Subjective Collaboration Perception in Mixed Reality

@@ Abstract
Mixed Reality (MR) enables rich, embodied collaboration; however, it is uncertain whether sensor- and system-logged behavioral signals capture how users experience that collaboration. This disconnect stems from a fundamental gap. Behavioral signals are observable and continuous, while collaboration is interpreted subjectively and shaped by internal states like presence, cognitive availability, and social awareness. Our core insight is that sensor signals serve as observable manifestations of subjective experiences in MR collaboration, and they can be captured through sensor data such as shared gaze, speech, spatial movement, and other system-logged performance metrics. We propose the Sensor-to-Subjective (S2S) Mapping Framework, a conceptual model that links observable interaction patterns to users’ subjective perceptions of collaboration and internal cognitive states through sensor-based indicators and task performance metrics. To evaluate this model, we conducted an exploratory study with 48 participants across 12 MR groups engaged in a collaborative image-sorting task. Our findings show a correlation between sensed behavior and perceived collaboration, particularly through shared attention and proximity.
@@
Modern MR systems generate extensive behavioral data through system logs and embedded sensors, capturing where users look [48], how they move [34], and when they speak [53]; however, these signals often provide only limited insights into users’ internal states [77]. Prior work shows that speech, gaze, and turn-taking are tightly connected to the internal perception of collaboration [11]. However, two core limitations remain: (1) it is unclear whether sensor data meaningfully reflects how participants perceive collaboration, or if key experiential dimensions go undetected. (2) Existing studies demonstrate specific bi-modal correlation mappings, such as gaze to presence [4], and physiological synchrony to co-presence [9], but stop short of offering a holistic objective-to-subjective mapping of collaboration. Bridging this gap is necessary to move from passive activity logging to human-aware MR systems that can infer and support collaboration quality in real time without relying on intrusive self-reports.
We address these gaps through a systematic, multimodal Sensor-to-Subjective (S2S) mapping framework, applied in a multi-perspective study of collaborative MR involving 48 participants across 12 groups. First, we analyze sensor-derived indicators of group interaction, including gaze-based shared attention, conversation dynamics through audio, and spatial proximity, which can be captured through embedded sensing in MR headsets [77, 73]. Second, we examine task performance using system logs, measuring timing, interaction patterns, and decision changes throughout the collaboration. Finally, we assess subjective experience through post-exposure questionnaires that capture participants’ perceptions of group dynamics as well as their individual sense of presence and cognitive effort. Together, these perspectives allow us to ask: Do observable group interaction patterns, as captured by MR headsets, reflect how participants experience collaboration? As an exploratory study, our objective is to identify promising alignments and generate hypotheses for future confirmatory research, rather than establishing a final predictive model for all MR contexts. We formalize this inquiry through three research questions:
@@
Do sensor-based indicators align with perceived subjective collaboration in MR?
@@
What do objective sensing and individual subjective experience reveal about group behavior in MR?
@@ 1.1 Group Behavior in Immersive Environments
Understanding collaborative group behavior has evolved from traditional settings to MR environments, where users interact in novel ways across co-located, remote, and hybrid configurations. Group behavior, is defined as all forms of interaction and activity within a group [22], and collaboration, the intentional, coordinated effort among members to work together toward shared goals [22], takes distinctive forms in MR contexts due to unique factors including free movement [45], embodied gesture [27], co-presence [56], and shared spatial context [44]. Research has examined MR collaboration through interface design [7], task coordination [75], communication patterns [3], and trust development [6], but it generally relies on external observations or outcomes rather than internal experiences of the participants. This creates a critical knowledge gap on how participants’ internal experiences align with the observable behaviors logged by MR systems.
@@ 1.2 Sensor-Based Indicators of Group Behavior
Sensor-based methods across ubiquitous and immersive computing characterize group behavior through multiple data streams (gaze, voice, motion, location) without relying on external observation or self-reporting. Remote collaboration systems, such as TeamSense [77] and CoCo [51] monitor nonverbal cues to track group cohesion, while wearable systems assess synchrony and proximity patterns during collaborative tasks [59]. In MR, headset-embedded sensors can track gaze, object manipulation, and gesture alignment, which researchers have shown correspond to collaboration quality [73, 26]. Studies have correlated gaze and gesture sharing with co-presence [4], and linked physiological synchrony and gaze with perceived collaboration quality [9]. Other systems enhance spatial awareness by embedding gaze cues [28], while multimodal integrations of gaze and gesture have been used to improve performance and presence [5]. Additional work shows high-accuracy cognitive load prediction using multimodal sensing [25] and demonstrates that even minimal AR cues can influence team experience [62]. Conversation-focused research further emphasizes how speech alignment, gesture, gaze, and turn-taking synchronize to create effective collaboration in co-located settings [61, 11]. These findings emphasize that collaboration is fundamentally multimodal, involving a convergence of signals across time.
Despite these insights, existing work remains limited in two ways: (1) Bi-modal Isolationism: isolated pairwise mappings (such as gaze to presence or physiological synchrony to co-presence) rather than producing a unifying mapping model and treating these as independent variables, and (2) Outcome-Centricity: lens-like views often aim to enhance either performance outcomes or structural interaction properties (such as coordination patterns and role distribution) without examining their relationship to participants’ subjective experiences. Our approach differs not simply by aggregating these modalities, but by shifting the focus from pairwise correlations to triadic alignment.
Prior work has explicitly recognized the need for structured frameworks that integrate multiple sensing modalities and interpret internal collaboration states. Barrett et al. call for multi-method frameworks that integrate sensor modalities to capture internal collaboration states [49], while social signal processing emphasizes the need to fuse multimodal behavioral signals to capture fine-grained group dynamics [62]. Systematic reviews also note the absence of frameworks that link objective data to users’ subjective experiences [17]. Despite these advances, the critical question of whether sensor-based behavioral indicators meaningfully reflect users’ internal perceptions of collaboration remains largely unexplored, particularly in co-located MR tasks using commercial headsets. Unlike a collection of bi-modal studies that treat sensors as stand-alone predictors, in this work, we position sensors, performance, and perception as interdependent nodes that validate one another through a unified interpretive lens.
@@ 1.3 Subjective Indicators of Group Behavior
Subjective self-report surveys remain central to understanding collaborative experiences. Standardized tools, such as post-task surveys, presence questionnaires [70, 52], and cognitive load surveys [23], and team cohesion surveys [77] gauge how users internalize virtual social presence [46], mutual understanding [76], and contribution equity [68], while collaborative design work leverages them to assess idea sharing and creative synergy [42]. Despite these rich insights, subjective measures are rarely linked systematically to sensor-based behavioral data. We address this disconnect by investigating whether MR-embedded signals (gaze, speech, proximity) meaningfully reflect these internal states, positioning this alignment as a core concern for human-centered sensing.
@@
2 The Sensor-to-Subjective (S2S) Mapping Framework
We propose the Sensor-to-Subjective (S2S) mapping framework to structure our investigation (Figure 1). S2S is an interpretive alignment model and theoretical structure designed to explain how high-frequency behavioral data relate to users’ internal collaborative experiences. The primary motivation for this alignment is the transition from passive logging to exploring active intervention. Rather than predicting subjective states from sensor data, we test whether and where system-observable sensor-derived interaction traces serve as an interpretive anchor for how collaboration is experienced. With S2S, we conceptualize collaborative group behavior across three interconnected perspectives (nodes): user-reported post-task reflections (subjective perceptions), observable system-level interaction traces (sensor-based indicators), and emergent outcomes from interaction (task performance metrics). We provide a foundation for MR systems to monitor, evaluate, and eventually enhance group dynamics autonomously. The critical distinction between S2S and existing bi-modal approaches that assume one view determines another [4, 9, 49, 61], is the shift from linear prediction to triadic synchronization. While prior work might ask ’Does gaze predict performance?’, we ask ’Does the sensor-performance alignment match the user’s subjective reality?’.
2.1 Framework Nodes: Indicators and Measures
2.1.1 Subjective Node: Perceptions of Collaboration
We define collaboration as not just joint actions, but how individuals interpret, feel, and reflect on their shared experience. In MR, social cues, task progress, and attention are distributed across a blend of physical and virtual elements, shaped by both individual states and group-level awareness. We treat this as the subjective perception of collaboration and operationalize it through individual-level and group-level dimensions.
Individual-Level Experience. Each user’s interpretation of collaboration is filtered through their sense of presence [55], cognitive availability [57], and personal contribution [14]. These passive background states directly influence how users perceive and make sense of social interaction [46, 74, 69]. At its most foundational, collaboration begins with how individuals experience being part of a group in each moment. In immersive settings, this includes:
Cognitive Load: The mental effort to manage task demands affects users’ ability to perceive coordination signals. A taxed user may miss coordination cues, while a cognitively available one may detect attention shifts or conversational breaks [15]. Thus, we hypothesize that cognitive load influences not only task performance [30] but also perceived collaboration quality [60].
Contribution Awareness: Users’ sense of how meaningfully others contributed reflects perceived participation balance [13]. This acts as a bridge between personal effort and social perception. It links internal judgment with observed peer input and becomes especially salient in unstructured, role-free MR settings.
These experiences are not isolated metrics but are shaped by real-time interpretations of others’ behaviors, and together they form the perceptual and cognitive basis for group-level reflections.
Group-Level Reflection. MR system tracks users individually, while users form post-task judgments about shared experience [58], such as whether they attended to the same objects, communicated effectively, and collaborated cohesively. We capture this through four constructs:
@@ 2.1.1 Subjective Node: Perceptions of Collaboration
Group Collaboration measures users’ overall sense of group effectiveness based on observed behaviors, internal judgments, and implicit comparison to collaborative expectations [19].
These individually collected measures reflect each user’s interpretation of the group, shaped by their state, engagement, and task involvement, forming subjective readings of a social process.
2.1.2 Objective Node: Sensor-Based Indicators
While subjective reports capture how collaboration feels, immersive systems log how it unfolds. Although these are individual-level measurements, they are not isolated; they reflect what each user does in relation to others in the group. The same user performing the task alone would not produce the same sensor traces [25]. These sensor-based signals (derived from sensor data and system instrumentation) are shaped by interaction: who a user looks at, how close they are to others, and when they speak or remain silent. Interpreting such distributed traces in group contexts is key to understanding collaboration objectively. We focus on three core indicators supported by prior work [73, 77] and accessible on commodity MR headsets [37, 2, 1]:
Conversation and speech activity have long been used to study group behavior across psychology, anthropology, and HCI [33, 12]. Recent work using ubiquitous sensors underscores the continued importance of conversation [32]. In MR, where interaction blends physical and virtual contexts, speech activity offers valuable insight into collaboration [73]. To this end, we extract speaking time, turn count, and participation variance (non-transcribing content) via the headset microphone. These metrics, known to reflect dominance and fluency [43], offer insight into how evenly members contribute with privacy preservation.
Shared attention, the ability to jointly focus on an object underpins coordination and social connection [71]. It has been widely studied in collaborative systems [8, 40], often through gaze-based detection[76]. We build on this by using MR headset eye-tracking to detect overlapping gaze on the same virtual object, enabling continuous, time-resolved logging of shared attention [40]. Such alignment reflects group awareness and may influence perceptions of collaboration, even without explicit communication. We aim to assess whether sensed shared attention reflects users’ perception of collaboration.
@@ 2.1.2 Objective Node: Sensor-Based Indicators
We focus on group-level distributions of proximity and conversation (overall closeness or speech participation) rather than directional relationships (who was closest to whom or who spoke to whom). Our goal is to capture collective engagement, not dominant roles. Prior work shows that proximity itself (regardless of direction) signals social engagement [16], and evenly distributed proximity and turn-taking indicate group cohesion [77]. We focus on subjective perceived group collaboration that is more likely influenced by general closeness than by specific spatial arrangements.
2.1.3 Performance Node: Emergent Task Outcomes
Beyond individual sensor signals and subjective reports, the S2S framework incorporates task-relevant interaction logs to measure the emergent outcomes of collaboration. These metrics serve as objective benchmarks for group efficiency and coordination complexity. We track three primary indicators (task details in §3.3.1):
Interaction balance and complexity log how often users interact with virtual content, the frequency of placement revisions (label changes), and the number of unique configurations explored before consensus. We focus on interaction balance rather than raw activity volume because participation equity reflects group cohesion and collective intelligence [72]. High activity can come from a single dominant user, while balanced interaction indicates distributed coordination toward shared goals. High variance in these interactions can reflect participation inequity or negotiation difficulty.
Coordination fluency monitors label overrides, which occur when one participant modifies a peer’s previous decision. This serves as a proxy for group negotiation and the alignment of shared mental models.
Temporal efficiency captured by task completion time is recorded as a coarse but vital measure of overall collaborative success.
These metrics allow us to evaluate whether specific behavioral data (participation balance, coordination, and negotiation of decisions) lead to more coherent and faster group outcomes.
2.2 Individual Reflections as Interpretive Anchors
@@
A central question in our exploratory study is how individual perceptions of collaboration relate to sensed behavioral traces. Rather than averaging subjective reports into a “single group state”, we examine the distribution of perceptions across the group to identify alignment or divergence. Do some users feel disengaged or overloaded? These differences are not treated as noise; they reflect varied interpretations of what the collaboration felt like to the user, such as whether they felt disengaged or cognitively overloaded. By linking subjective internal perceptions (§2.1.1) with sensor-based indicators (§2.1.2) and task performance metrics (§2.1.3), we ask: Do shared gaze events map onto users’ awareness of shared attention? Does proximity reflect perceived cohesion? Do balanced interactions reflect users’ perception of their own and peers’ meaningful contributions to the task?
@@
Importantly, in S2S, subjective states are not treated as downstream outcomes but as interpretive anchors that ground how sensor traces and task outcomes are understood; independent but complementary accounts of collaboration. This allows us to understand what groups did and how that behavior was experienced, yielding insights into how systems might better reflect, support, or adapt to group behavior in MR environments in the future. Together, these indicators form a continuous stream of behavioral data, interpreted as isolated features but also as observable patterns of interaction that may or may not align with users’ subjective interpretations of collaboration, avoiding the false precision that arises when sensors are interpreted in isolation. Rather than labeling groups as “effective” or “cohesive,” our goal is to assess whether and how the system’s observations reflect how collaboration was experienced.
2.3 Functional Relational Pathways
The substance of the S2S framework lies in the functional mappings between its nodes, which allow us to explore the alignment between behavioral data and human experience. We define three primary pathways that guide the structure of our investigation:
2.3.1 Reflective Interaction Mapping (Sensor Subj.)